# Few-shot Class-incremental Learning: A Survey

### v2

@@ 4.4.1 Performance Comparison
In this section, we summarize the performance of mainstream FSCIC methods. Since not all relevant methods are open-source, and there are certain differences in their implementation conditions and configurations (such as different backbones, the fusion with other features, and different learning paradigms), it is impractical to compare all FSCIC methods in a unified and fair manner. Therefore, we summarize the performance of FSCIC methods on three commonly used benchmark datasets in Tab. III, including miniImageNet, CIFAR-100, and CUB-200. To fully demonstrate the characteristics of each method, Tab. III provides their types and specific taxonomy categories. In addition, we provide the backbone used by each method in this table. As some methods introduce additional auxiliary factors, we have specially set an ”extra” column in the table to summarize the additional auxiliary factors introduced by each method. The performance of FSCIC methods is primarily evaluated by measuring the accuracy achieved on different incremental sessions, AA across all sessions, and PD values. Given the space limitations, we only provide accuracy for the starting and ending sessions, AA, and PD. Moreover, we summarize the highlights of each method in this table.
TABLE III: The performance of FSCIC methods. The accuracy obtained on the start and end sessions, self-supervised, and unlabeled data are abbreviated to SA, EA, SS, and un-data, respectively. (In %)
Type Taxonomy Methods Venue Backbone (ResNet) Extra Factor miniImageNet CIFAR-100 CUB-200 Highlights
SA EA AA PD↓ SA EA AA PD↓ SA EA AA PD↓
Data-based Data Replay FSIL-GAN [68] ACM MM 18 - 69.87 46.14 56.40 23.73 70.14 46.61 55.31 23.53 81.07 59.13 69.26 21.94
@@
Data Replay
ERDFR [29] ECCV 18 & 20 - 71.84 48.21 58.02 23.63 74.40 50.14 60.78 24.26 75.90 52.39 61.52 23.51
Proved the effectiveness of data replay in FSCIL and proposed a data-free replay method
@@
FDD [44] PRAI 18 - 64.14 42.01 52.44 22.13 64.94 41.73 52.52 23.21 - - - -
Fix the shallow layers and fine-tune the deep layers with KD loss and cross-entropy loss
Pseudo Scenarios SPPR [69] CVPR 18 - 61.45 41.92 52.76 19.53 63.97 43.32 54.45 20.65 68.68 37.33 49.32 31.35
@@
Pseudo Scenarios FACT [33] CVPR 18 & 20 - 72.56 50.49 60.70 22.07 74.60 52.10 62.24 22.50 75.90 56.94 64.42 18.96
@@
ALICE [47] ECCV 18 - 80.60 55.70 63.99 24.90 79.00 54.10 63.21 24.90 77.40 60.10 65.75 17.30
ALICE, which uses angular penalty loss for discriminated and generalized feature learning
Structure-based Dynamic Structure TOPIC [13] CVPR 18 - 61.31 24.42 39.64 36.89 64.10 29.37 42.62 34.73 68.68 26.28 43.92 42.40
@@
CEC [14] CVPR 18 & 20 - 72.00 47.63 57.75 24.37 73.07 49.14 59.53 23.93 75.85 52.28 61.33 23.57
Propose CEC framework based on GAT to propagate context information, and it was trained pseudo-incremental sessions
Representation Learning
Dynamic Structure
FeSSSS [72] CVPR 18 & 20
Extra SS
module
81.50 58.87 68.24 22.63 75.35 50.23 60.92 25.12 79.60 52.98 62.86 26.62
Explored self-supervised learning in FSCIL and proposed a novel fusion strategy for supervised and self-supervised features
Dynamic Structure DSN [35] TPAMI 18 - 66.95 45.89 54.39 21.06 73.00 50.00 60.14 23.00 80.86 63.21 71.02 17.65
@@
Optimization-based Representation Learning SFbFSCIL [79] ICCV 18 - 61.37 42.23 50.73 19.14 62.00 41.69 50.86 20.31 68.78 43.23 51.84 25.55
@@
Representation Learning
FSLL+SS [31] AAAI 18 - 68.85 43.92 53.92 24.93 66.76 39.57 48.71 27.19 75.63 55.82 62.62 19.81
FSLL trains only a few selected parameters to limit overfitting, and leverages self-supervision
Representation Learning
Attention
C-FSCIL [34] CVPR 12 - 76.40 51.41 61.61 24.99 77.47 50.47 61.64 27.00 - - - -
Proposed C-FSCIL based on a meta-learned feature extractor, fully connected layer, and dynamic memory
Representation Learning CFSCIL [85] CVPRW 18
Extra SS
module
81.50 58.87 68.24 22.63 - - - - 79.60 52.98 62.86 26.62
Proposed variable FSCIL for the first time and investigated the use of self-supervised learning for it
Representation Learning MgSvF [32] TPAMI 18 - 63.09 45.76 52.87 17.33 74.24 51.40 61.67 22.84 72.29 54.33 62.37 17.96
MgSvF based on frequency-aware regularization and feature space composition to balance old and new knowledge
Representation Learning S3C [86] ECCV 18 & 20
Extra SS
module
76.89 52.37 62.77 24.52 78.66 53.99 64.90 24.67 80.62 58.95 67.03 21.67
S3C integrated stochastic classifiers with self-supervised learning, and FSCIL-im and FSCIL-lb settings were proposed
Representation Learning CLOM [36] NIPS 18 & 20 - 73.08 48.00 58.48 25.08 74.20 50.25 60.57 23.95 79.57 59.58 67.17 19.99
@@
Representation Learning WaRP [80] ICLR 18 & 20 - 72.99 50.65 59.69 22.34 80.31 54.74 65.82 25.57 77.74 57.01 64.66 20.73
WaRP, a weight space rotation process that compresses old knowledge into key parameters, allowing fine-tuning without forgetting
Representation Learning
NC-FSCIL [82] ICLR 12 & 18 - 84.02 58.31 67.82 25.71 82.52 56.11 67.50 26.41 80.45 59.44 67.28 21.01
Proposed a neural collapse inspired framework
@@
SaKD [27] CVPR 18 - 61.33 38.73 48.20 22.60 64.03 34.94 45.53 29.09 68.23 32.96 46.13 35.27
Proposed a semantic-aware distillation method and an attention-driven alignment strategy to mitigate catastrophic forgetting
Knowledge Distillation SSFSCIL [93] ICIP 18
Extra
un-data
62.99 41.23 51.24 21.76 64.09 44.69 54.53 19.40 69.89 35.25 50.78 34.64
Introduced semi-supervised FSCIL setting for the first time, and provided comprehensive benchmarks
Knowledge Distillation ERL++ [92] AAAI 18 - 61.71 40.77 49.84 20.94 73.70 48.25 59.31 25.45 73.52 52.28 61.18 21.24
@@
Knowledge Distillation Us-KD [45] TMM 18
Extra
un-data
72.35 50.47 58.79 21.88 76.85 54.42 61.80 22.43 74.69 55.54 62.89 19.15
Incorporated KD and uncertainty-guided semi-supervised learning to perform FSCIL
Knowledge Distillation SSFE-Net [64] WACV 18
Extra SS
module
72.06 47.73 57.74 24.33 - - - - 76.38 54.28 62.99 22.10
Proposed a self-supervised learning and KD framework to enhance feature extraction and new datasets
@@
CaBD [73] CVPR 18 - 74.65 53.39 62.06 21.26 79.45 56.56 66.55 22.89 79.12 63.81 68.61 15.31
Proposed a class-aware bilateral distillation method by adaptively drawing knowledge from two complementary teachers
Knowledge Distillation UaD-CE [46] TNNLS 18
Extra
un-data
72.35 50.52 58.82 21.83 75.55 54.50 63.93 21.05 75.17 60.72 65.70 14.45
Proposing UaD-CE for semi-supervised FSCIL with uncertainty-guided distillation and class equilibrium
@@
Meta Learning CSR [95] ICDMW 18 - 67.67 44.52 54.11 23.15 72.02 49.00 59.07 23.02 74.69 55.09 62.32 19.60
@@
LIMIT [16] TPAMI 18 & 20 - 72.32 49.19 59.06 23.13 73.81 51.23 61.84 22.58 75.89 57.41 65.48 18.48
LIMIT, a meta-learning based paradigm, which synthesizes fake tasks to build a generalized feature space for unseen tasks
Others F2M [30] NIPS 18 - 67.28 44.65 54.89 22.63 64.71 44.67 53.65 20.04 81.07 60.26 69.49 20.81
@@
For FSCIC methods, the performance of the backbone achieved on the base session is crucial for subsequent IL. Through the analysis of the SA of the three datasets in Tab III, it can be found that the top five performance methods on miniImagNet are: NC-FSCIL [82] with 84.02%, FeSSSS [72] and CFSCIL [85] with 81.50%, ALICE [47] with 80.60%, and S3C [86] with 76.89%. On CIFAR-100, the top five methods are: NC-FSCIL [82] with 82.52%, WaRP [80] with 80.31%, CaBD [73] with 79.45%, ALICE with 79.00%, and S3C with 78.66%. The top five methods on the CUB-200 dataset are F2M [30] and FSIL-GAN [68] with 81.07%, DSN [35] with 80.86%, S3C with 80.62%, and NC-FSCIL with 80.45%. It is noticeable that the NC-FSCIL based on the neural collapse theory and the S3C that integrates the self-supervised learning features have achieved excellent performance on all three benchmark datasets. The ALICE framework that utilizes metric learning and pseudo-data synthesis has performed well on miniImagNet and CIFAR-100. Furthermore, combining the performance of other methods on the three datasets, it can be found that introducing self-supervised learning features can help the FSCIC model achieve good performance on the base session. The appropriate metric learning method and neural collapse theory can also improve the model’s performance on the base session. Apart from that, it has been found that almost all methods that achieved top five performance on the base session also achieved top five performance in terms of the AA index. It reflects the influence of the performance achieved on the base session.
@@
The performance obtained in the final session reflects the learning capability of the FSCIC model for incremental classes and the stability of keeping old knowledge. However, as the model learned in each incremental session will be tested on all seen classes, and the number of classes involved in the base session is large, the performance obtained on each session cannot fully represent the model’s IL ability. In contrast, the PD value can better reflect the model’s ability to resist forgetting. In Tab. III, the top five methods with the lowest PD values on miniImageNet are: MgSvF [32] with 17.33%, SFbFSCIL [79] with 19.14%, SPPR [69] with 19.53%, ERL++ [92] with 20.94%, and DSN [35] with 21.06%. The top five methods on CIFAR-100 are: SSFSCIL [93] with 19.40%, F2M [30] with 20.04%, SFbFSCIL with 20.31%, SPPR with 20.65%, and UaD-CE [46] with 21.05%. The top five methods on CUB-200 are: UaD-CE with 14.45%, CaBD [73] with 15.31%, ALICE [47] with 17.30%, DSN with 17.65%, and MgSvF with 17.96%. It can be seen that SPPR, which constructs pseudo incremental sessions, and SFbFSCIL, which is based on feature space fusion and VAE feature synthesis, both have achieved good PD values on the first two datasets. UaD-CE, based on KD and semi-supervised learning, also achieved good PD values on the last two datasets. Based on dynamic network structure, DSN and MgSvF, based on frequency domain analysis, have shown good performance on the first and last datasets. Furthermore, combining the performance of other methods on the three datasets, it can be found that techniques such as KD, pseudo-incremental scenario construction, dynamic structures, and feature optimization can effectively alleviate the catastrophic forgetting problem.
@@ 5.3.1 Performance Comparison
In this section, we summarize the performance of FSCIOD methods. We summarize the performance of relevant methods on COCO and VOC in Tab. IV. To fully elucidate the attributes of each method, Tab. IV includes their types and specific taxonomy categories. In addition, the backbone employed by each method is furnished in this table. Because some methods can achieve object detection and instance segmentation simultaneously, we have added a ”task” column in Tab. IV to denote performance on related tasks. There are two evaluation strategies for FSCIOD methods: normal evaluation on COCO and cross-dataset evaluation on COCO and VOC. Given that relevant works adopt two combinations of evaluation metrics: AP and AR, and AP and AP50, we provide AP, AP50, and AR values in Tab. IV. Furthermore, we distill the highlights of each method in Tab IV.
TABLE IV: The performance of FSCIOD methods. The taxonomy is abbreviated to taxo. (In %)
@@
AP AP50 AR AP AP50 AR AP AP50 AR AP AP50 AR
Anchor-free CentreNet-based ONCE [62] CVPR 50 D 1 17.90 - 19.50 0.70 - 6.30 13.60 - 16.20 - - - Proposed FSCIOD setting and introduced the first work, ONCE
@@
SS [99] TCSVT 50 D 1 26.90 - 25.80 0.90 - 4.20 20.40 - 20.40 1.50 2.30 6.10 Proposed new models, redesigning the CenterNet and incorporating a novel meta-learning method, MAML, to perform FSCIOD task
@@
FCOS-based Sylph [97] CVPR 50 D 1 37.60 - - 1.10 - - 28.48 - - - - - Introduced FCOS-based Sylph, decoupling object detection into classification and localization
@@
MCH [100] PRL 50 D 1 36.90 - - 0.40 - - 27.70 - - 1.00 - - Introduced MCH and BPMCH, human memory-inspired models, outperforming ONCE by effectively transferring knowledge from base to novel classes
@@
DETR -based Incremental -DETR [98] AAAI 50 D 1 29.40 47.10 - 1.90 2.70 - 22.50 36.00 - 4.10 6.60 - Developed Incremental-DETA for FSCID, which uses self-supervised learning and a fine-tuning strategy
@@
Anchor-based Mask RCNN-based iMTFA [96] CVPR 50 D 1 27.81 40.11 - 3.23 5.89 - 21.67 31.55 - - - - Proposed instance segmentation setting in FSCIL and introduced the first work, iMTFA, which can perform both instance segmentation and object detection
@@
iFS-RCNN [107] CVPR 50 D 1 40.08 - - 4.54 - - 31.19 - - - - - Introduced iFS-RCNN, an extension of Mask-RCNN, leveraging probit function and uncertainty-guided bounding box prediction for instance segmentation and object detection in FSCIL
@@
Given that the FSCIOD evaluation is usually conducted under different sample shots, we analyze and summarize based on the overall performance of relevant methods. It can be found from Tab. IV, the top three performance methods for object detection achieved on base classes are Sylph [97], iFS-RCNN [107], and MCH [100]. The top three performance methods for novel COCO classes are iFS-RCNN, Incremental-DETR [98], and iMTFA [96]. The top three methods for overall performance on COCO are iFS-RCNN, Sylph, and MCH. Among all methods for cross-dataset evaluation on VOC, the top three performers are: Incremental-DETR, BPMCH [100], and MCH. Therefore, it can be seen that iFS-RCNN based on complex Mask RCNN yields the best results, and Sylph and MCH, which are based on simple FCOS, also show good performance. In instance segmentation, only anchor-based methods have conducted the relevant evaluation, among which iMTFA has the overall best incremental segmentation ability on novel classes, but iFS-RCNN performs best on base classes. In summary, anchor-based methods are suitable for object detection and instance segmentation scenarios, with excellent performance but more complex structures; anchor-free methods are suitable for application scenarios requiring lower framework complexity and can achieve performance slightly inferior to anchor-based methods.

### v3

@@ 4.4.1 Performance Comparison
In this section, we summarize the performance of mainstream FSCIC methods. Since not all relevant methods are open-source and their implementation conditions and configurations (such as different backbone networks, feature fusion with other methods, and different learning paradigms) vary, we summarize the performance of mainstream FSCIC methods with similar backbones on three commonly used benchmark datasets in Table IV, including miniImageNet, CIFAR-100, and CUB-200, to enable a fair comparison as much as possible. The performance values are extracted from corresponding papers. To fully demonstrate the characteristics of each method, Tab. IV provides their types and specific taxonomy categories. In addition, we provide the backbone used by each method in this table. For methods with too many additional factors, we provide a supplementary table in the appendix for reference. Since some methods introduce extra auxiliary factors, we have specially included an ”extra factor” column in the table to summarize these factors for each method. The performance of FSCIC methods is primarily evaluated by measuring the accuracy achieved on different incremental sessions, AA across all sessions, and PD values. Given the space limitations, we only provide accuracy for the starting and ending sessions, AA, and PD. Moreover, we summarize the highlights of each method in these tables.
TABLE IV: The performance of mainstream FSCIC methods. The data are extracted from the original papers. For miniImageNet and CUB-200, all the methods employ ResNet18 as the backbone. For CIFAR-100, related methods have two different settings, ResNet18 and ResNet20. To provide a comprehensive comparison, we report them in this table together, with the details noted. The accuracy of the first and last sessions is abbreviated as SA and EA. We use “-” to mark the dataset without reporting in the original papers. The best results are bold and underlined, while the second-best are underlined only. (In %)
Type Taxonomy Methods Venue miniImageNet CIFAR-100 CUB-200 Highlights
Backbone
SA EA AA PD↓
Backbone
SA EA AA PD↓
Backbone
SA EA AA PD↓
Data-based Data Replay FSIL-GAN [68]
ACM MM 22
18 69.87 46.14 56.40 23.73 18 70.14 46.61 55.31 23.53 18 81.07 59.13 69.26 21.94
@@
Data Replay ERDFR [19] ECCV 22 18 71.84 48.21 58.02 23.63 20 74.40 50.14 60.78 24.26 18 75.90 52.39 61.52 23.51
Proved the effectiveness of DR in FSCIL and proposed a data-free replay method
@@
FDD [44] PRAI 22 18 64.14 42.01 52.44 22.13 18 64.94 41.73 52.52 23.21 18 - - - -
Fixed the shallow layers and fine-tune the deep layers with KD loss and CE loss
Pseudo Scenarios SPPR [69] CVPR 21 18 61.45 41.92 52.76 19.53 18 63.97 43.32 54.45 20.65 18 68.68 37.33 49.32 31.35
@@
Pseudo Scenarios FACT [23] CVPR 22 18 72.56 50.49 60.70 22.07 20 74.60 52.10 62.24 22.50 18 75.90 56.94 64.42 18.96
@@
ALICE [47] ECCV 22 18 80.60 55.70 63.99 24.90 18 79.00 54.10 63.21 24.90 18 77.40 60.10 65.75 17.30
ALICE used angular penalty loss for discriminated and generalized feature learning
Pseudo Scenarios
Representation Learning
SAVC [96] CVPR 23 18 81.12 57.11 67.05 24.01 20 78.77 53.12 63.63 25.65 18 81.85 62.50 69.35 19.35
Used supervised contrast learning and virtual classes to initialize backbone so that it can have good generalization performance
Structure-based Dynamic Structure TOPIC [13] CVPR 20 18 61.31 24.42 39.64 36.89 18 64.10 29.37 42.62 34.73 18 68.68 26.28 43.92 42.40
@@
CEC [14] CVPR 21 18 72.00 47.63 57.75 24.37 20 73.07 49.14 59.53 23.93 18 75.85 52.28 61.33 23.57
Proposed CEC framework based on GAT to propagate context information, and it was trained pseudo-incremental sessions
Dynamic Structure DSN [25] TPAMI 22 18 66.95 45.89 54.39 21.06 18 73.00 50.00 60.14 23.00 18 80.86 63.21 71.02 17.65
@@
Optimization-based Representation Learning SFbFSCIL [79] ICCV 21 18 61.37 42.23 50.73 19.14 18 62.00 41.69 50.86 20.31 18 68.78 43.23 51.84 25.55
@@
Representation Learning FSLL+SS [21] AAAI 21 18 68.85 43.92 53.92 24.93 18 66.76 39.57 48.71 27.19 18 75.63 55.82 62.62 19.81
FSLL trained only a few selected parameters to limit overfitting, and leverages self-supervision
Representation Learning MgSvF [22] TPAMI 21 18 63.09 45.76 52.87 17.33 18 74.24 51.40 61.67 22.84 18 72.29 54.33 62.37 17.96
MgSvF used frequency-aware regularization and feature space composition to balance old and new knowledge
Representation Learning CLOM [26] NeurIPS 22 18 73.08 48.00 58.48 25.08 20 74.20 50.25 60.57 23.95 18 79.57 59.58 67.17 19.99
@@
Representation Learning WaRP [80] ICLR 23 18 72.99 50.65 59.69 22.34 20 80.31 54.74 65.82 25.57 18 77.74 57.01 64.66 20.73
WaRP, a weight space rotation process that compressed old knowledge into key parameters, allowing fine-tuning without forgetting
Representation Learning TEEN [97] NeurIPS 23 18 73.53 52.08 61.45 21.45 20 74.92 52.64 63.10 22.28 18 77.26 59.31 66.63 18.13
TEEN, a training-free calibration strategy, enhanced new class discriminability by fusing new and weighted base class prototypes.
@@
SaKD [17] CVPR 21 18 61.33 38.73 48.20 22.60 18 64.03 34.94 45.53 29.09 18 68.23 32.96 46.13 35.27
Proposed a semantic-aware distillation method and an attention driven alignment strategy to mitigate catastrophic forgetting
Knowledge Distillation ERL++ [92] AAAI 21 18 61.71 40.77 49.84 20.94 18 73.70 48.25 59.31 25.45 18 73.52 52.28 61.18 21.24
@@
BiDistFSCIL [73] CVPR 23 18 74.65 52.22 61.42 22.43 18 79.45 55.88 66.14 23.57 18 79.12 60.93 67.34 18.19
It proposed a KD strategy with two teacher models, designed a two-branch network, and used an attention mechanism to aggregate the predictions from both branches.
@@
Meta Learning CSR [95] ICDMW 21 18 67.67 44.52 54.11 23.15 18 72.02 49.00 59.07 23.02 18 74.69 55.09 62.32 19.60
@@
LIMIT [16] TPAMI 22 18 72.32 49.19 59.06 23.13 20 73.81 51.23 61.84 22.58 18 75.89 57.41 65.48 18.48
LIMIT, a meta-learning based paradigm, which synthesized fake tasks to build a generalizable feature space for unseen tasks
Others F2M [20] NeurIPS 21 18 67.28 44.65 54.89 22.63 18 64.71 44.67 53.65 20.04 18 81.07 60.26 69.49 20.81
@@
For FSCIC methods, the performance of the backbone achieved on the base session is crucial for subsequent IL. From the analysis of SA across the three datasets in Tab. IV, it can be seen that the top five methods on miniImageNet are: SAVC [96] (81.12%), ALICE [47] (80.60%), BiDistFSCIL [73] (74.65%), TEEN [97] (73.53%), and CLOM [26] (73.08%). On CIFAR-100, the top five methods are: WaRP [80] (80.31%), BiDistFSCIL (79.45%), ALICE (79.00%), SAVC (78.77%), and TEEN (74.92%). On the CUB-200 dataset, the top five methods are: SAVC (81.85%), F2M [20] and FSIL-GAN [68] (81.07%), DSN [25] (80.86%), and BiDistFSCIL (79.12%). It is noteworthy that the SAVC, based on virtual class synthesis, achieved the best initial performance on miniImageNet and CUB-200; the ALICE framework, which leverages metric learning and pseudo-data synthesis, also performed exceptionally well on these datasets. This indicates that the virtual class strategy is highly effective in improving performance in the initial session. Apart from that, it has been found that almost all methods that achieved top five performance on the base session also achieved top five performance in terms of the AA index. It reflects the influence of the performance achieved on the base session.
@@
The performance obtained in the final session reflects the learning capability of the FSCIC model for incremental classes and the stability of keeping old knowledge. However, as the model learned in each incremental session will be tested on all seen classes, and the number of classes involved in the base session is large, the performance obtained on each session cannot fully represent the model’s IL ability. In contrast, the PD value can better reflect the model’s ability to resist forgetting. In Tab. IV, the top five methods with the lowest PD values on miniImageNet are: MgSvF [22] (17.33%), SFbFSCIL [79] (19.14%), SPPR [69] (19.53%), ERL++ [92] (20.94%), and DSN (21.06%). On CIFAR-100, the top five methods are: F2M [20] (20.04%), SFbFSCIL (20.31%), SPPR (20.65%), TEEN (22.28%), and FACT [23] (22.50%). On CUB-200, the top five methods are: ALICE (17.30%), DSN (17.65%), MgSvF (17.96%), TEEN (18.13%), and BiDistFSCIL(18.19%). It can be seen that SPPR, which constructs pseudo incremental sessions, and SFbFSCIL, which is based on feature space fusion and VAE feature synthesis, both have achieved good PD values on the first two datasets. DSN, based on dynamic network structure, and MgSvF, based on frequency domain analysis, performed well on the first and last datasets. Furthermore, combining the performance of other methods on the three datasets, it can be found that techniques such as KD, pseudo-incremental scenario construction, dynamic structures, and feature optimization can effectively alleviate the catastrophic forgetting problem.
@@ 5.3.1 Performance Comparison
In this section, we summarize the performance of FSCIOD methods. We summarize the performance of relevant methods on COCO and VOC in Tab. V. To fully elucidate the attributes of each method, Tab. V includes their types and specific taxonomy categories. In addition, the backbone employed by each method is furnished in this table. Because some methods can achieve object detection and instance segmentation simultaneously, we have added a ”task” column in Tab. V to denote performance on related tasks. FSCIOD methods are evaluated in two ways: standard evaluation on COCO and cross-dataset evaluation on COCO and VOC. Tab. V presents mAP, mAP50, and mAR values, along with method highlights.
TABLE V: The performance of FSCIOD methods. The data are extracted from the original papers. The taxonomy is abbreviated to taxo. We use “-” to mark the dataset without reporting in the original papers. (In %)
@@
mAP mAP50 mAR mAP mAP50 mAR mAP mAP50 mAR mAP mAP50 mAR
Anchor-free CentreNet-based ONCE [62] CVPR 20 50 D 1 17.90 - 19.50 0.70 - 6.30 13.60 - 16.20 - - - Proposed FSCIOD setting and introduced the first work, ONCE
@@
SS [101] TCSVT 21 50 D 1 26.90 - 25.80 0.90 - 4.20 20.40 - 20.40 1.50 2.30 6.10 Proposed new models, redesigning the CenterNet and incorporating a novel meta-learning method, MAML, to perform FSCIOD task
@@
FCOS-based Sylph [99] CVPR 22 50 D 1 37.60 - - 1.10 - - 28.48 - - - - - Introduced FCOS-based Sylph, decoupling object detection into classification and localization
@@
MCH [102] PRL 22 50 D 1 36.90 - - 0.40 - - 27.70 - - 1.00 - - Introduced MCH and BPMCH, human memory-inspired models, outperforming ONCE by effectively transferring knowledge from base to novel classes
@@
DETR -based Incremental -DETR [100] AAAI 23 50 D 1 29.40 47.10 - 1.90 2.70 - 22.50 36.00 - 4.10 6.60 - Developed Incremental-DETR for FSCID, which uses self-supervised learning and a fine-tuning strategy
@@
Anchor-based Mask RCNN-based iMTFA [98] CVPR 21 50 D 1 27.81 40.11 - 3.23 5.89 - 21.67 31.55 - - - - Proposed instance segmentation setting in FSCIL and introduced the first work, iMTFA, which can perform both instance segmentation and object detection
@@
iFS-RCNN [109] CVPR 22 50 D 1 40.08 - - 4.54 - - 31.19 - - - - - Introduced iFS-RCNN, an extension of Mask-RCNN, leveraging probit function and uncertainty-guided bounding box prediction for instance segmentation and object detection in FSCIL
@@
Given that the FSCIOD evaluation is usually conducted under different sample shots, we analyze and summarize based on the overall performance of relevant methods. It can be found from Tab. V, the top three performance methods for object detection achieved on base classes are Sylph [99], iFS-RCNN [109], and MCH [102]. The top three performance methods for novel COCO classes are iFS-RCNN, Incremental-DETR [100], and iMTFA [98]. The top three methods for overall performance on COCO are iFS-RCNN, Sylph, and MCH. Among all methods for cross-dataset evaluation on VOC, the top three performers are: Incremental-DETR, BPMCH [102], and MCH. Therefore, it can be seen that iFS-RCNN based on complex Mask RCNN yields the best results, and Sylph and MCH, which are based on simple FCOS, also show good performance. In instance segmentation, only anchor-based methods have conducted the relevant evaluation, among which iMTFA has the overall best incremental segmentation ability on novel classes, but iFS-RCNN performs best on base classes. In summary, anchor-based methods are suitable for object detection and instance segmentation scenarios, with excellent performance but more complex structures; anchor-free methods are suitable for application scenarios requiring lower framework complexity and can achieve performance slightly inferior to anchor-based methods.